OpenStudy (anonymous):

solve 6(5-2z)+5z=8z-3(16-3z) please show work

OpenStudy (anonymous):

\[30-12z+5z=8z-48+9z\] \[30-7z=17z-48\] \[78=24z\] \[78/24=24z/24\] \[z=3.25\]

OpenStudy (anonymous):

multiply and you get 30-12z +5z=8z-48+9z move all the numbers with z to one side and numbers without z to the other 30+48=12z-5z+8z+9z remember the signs change when you move them from one side to another then combine! 78= 24z divide both sides by 24 z=3.25